mathslover (mathslover):

Prove that : \[\large{(x^3+y^3+z^3)^2 \ge \frac{9}{8}(x^2+yz)(y^2+zx)(z^2+xy)}\] where x,y and z are non - negative real numbers.

mathslover (mathslover):

should we expand here in RHS or in LHS ? I did that in RHS but that was too much complicated ...result

mathslover (mathslover):

@Callisto @myininaya @UnkleRhaukus

OpenStudy (unklerhaukus):

\[\text{if } \quad x=y=z=1\] \[9\geq\frac98(8)\]

mathslover (mathslover):

is it given that x = y = z = 1 ? @UnkleRhaukus

OpenStudy (unklerhaukus):

the equation is symmetric in the three variables i think x=y=z=1 is a important point on the graph

mathslover (mathslover):

Hmn but I don't think so as this as a correct way..

OpenStudy (anonymous):

I am still waiting for @mukushla here... He is good at inequalities..

OpenStudy (experimentx):

try proving this \[ \large{(x^3+y^3+z^3)^2 \ge \frac{3}{8}((x^2+yz)^3 +(y^2+zx)^3 + (z^2+xy)^3)}\]

OpenStudy (experimentx):

not sure though ...

mathslover (mathslover):

Hmn.... I will like to try that @experimentX . .... but please.... r u sure that this would help?

OpenStudy (experimentx):

might ... I'm extremely busy due to exams!!

mathslover (mathslover):

Oh k no problem.... best of luck for your exams....

mathslover (mathslover):

@TuringTest and @KingGeorge @amistre64 might like to view this when they will be online

OpenStudy (anonymous):

Here what I did so far which leads to the credibility that that the inequality is true: If \[ h(x,y,z)=\frac{\left(x^3+y^3+z^3\right)^2}{\left(x z+y^2\right) \left(x^2+y z\right) \left(x y+z^2\right)}\\ \] I computed its gradient. G and I verified that G(x,x,x)=0. Also \[ h(x,x,x)=\frac 98 \] One has to show that h has a minimum at those points and we will be done. I am not there.

mathslover (mathslover):

must say that I did the following and I hpe that I am very near to the answer

OpenStudy (anonymous):

emm dont feorget u missing 3abc\[a^3+b^3+c^3-3abc=(a+b+c)(a^2+b^2+c^2-ab-ac-bc)\]

OpenStudy (anonymous):

and of course \[xy+yz+zx\le x^2+y^2+z^2\]

mathslover (mathslover):

oh k I will try that again using your given hints.... @mukushla

OpenStudy (anonymous):

\[27(x^{2}+yz)(y^{2}+zx)(z^{2}+xy)\leq (x^{2}+y^{2}+z^{2}+xy+yz+zx)^{3}\] \(\leq 8(x^{2}+y^{2}+z^{2})^{3}\)

OpenStudy (anonymous):

But we need \( \left( x^3+y^3+ z^3 \right)^2 \) and not \( \left( x^2+y^2+ z^2 \right)^3 \) Am I missing something?

OpenStudy (anonymous):

im stuck with that sir : to proving\[(x^{2}+y^{2}+z^{2})^{3} \le3 (x^{3}+y^{3}+z^{3})^{2}\]

OpenStudy (anonymous):

i thought its obvious but now im stuck there

mathslover (mathslover):

no problem @mukushla actually @vishweshshrimali5 knew the correct answer and solution as he has just attempted that question today in olympiad and he says that he has a correct prooof.... He just challenged me... and I probably won the challenge... as I atleast tried it ..:P as per my opiniion that is all about " expansion " .... but still there will be suspense But today he is having rest I will let him know about this discussion and I am sure that he will declare the right proof here asap...

OpenStudy (anonymous):

Proving the above inequality is the same as proving the original one.

OpenStudy (anonymous):

A suggestion: The inequality's both sides are SYMMETRIC POLYNOMIALS in x,y,z And so they are expressible in products of lower order symmetric polynomials. Now lower order then 3 is either 2 or 1 and for these degrees a huge arsenal of harmonic, geometric and arithmetic MEAN inequalities is well known. (Mukusha probably knew the idea and he only showed the impressive consequence , not the internal principle)

OpenStudy (anonymous):

power mean inequality\[ \sqrt[3]{\frac{x^{3}+y^{3}+z^{3}}{3}}\geq\sqrt{\frac{x^{2}+y^{2}+z^{2}}{3}} \]so\[(x^{2}+y^{2}+z^{2})^{3} \le3 (x^{3}+y^{3}+z^{3})^{2}\]

OpenStudy (anonymous):

@mukushla Here is a challenging observation for an olympian: in every symmetric polynomial like here BOTH sides you can reduce the number of variables by 1 using uniform division by x^2y^2z^2 and intrducing 2 new ratio variables

mathslover (mathslover):

thanks guys...

OpenStudy (anonymous):

np :)